The police car is chasing the getaway car of bank robbers 7 1/2 km ahead. If the speed of the police car is 5 kph more than that of the getaway car, how long will it take the police car to overtake the getaway car?
Found 2 solutions by mananth, jrfrunner:
Answer by mananth(16949)

The police car is chasing the getaway car of bank robbers 7 1/2 km ahead. If the speed of the police car is 5 kph more than that of the getaway car, how long will it take the police car to overtake the getaway car?
...
catch up speed = 5 km/h
distance = 15/2 km
...
time they catch = (5)/(15/2)
=5*2/15
=2/3 hours or 40 minutes
...
m.ananth@hotmail.ca

Answer by jrfrunner(365)

d-distance to make up= 7.5 km
r=speed differential= 5kph
--
d=r*t or t=d/r
==
t=7.5/5=1.5 hours or 90 minutes